Find the equation of the line through (-1,2) that is perpendicular to 3x-y=4
Answer by checkley71(8403)

3x-y=4
-y=-3x+4
y=3x-4 thus the slope of this line is 3. A line perpendicular to this one has a slope of -1/3. Using this slope & the two points we have the line formula of--
y=mx+b where m=slope & b=y intercept.
2=(-1/3)*-1+b
2=1/3+b
b=2-1/3
b=1 2/3 or 5/3
Thus this line is y=-1/3x+5/3